OpenStudy (anonymous):

simplify the complex numbers: -10/5-7i How do you know to conjugate this problem because of the -7i in the denominator?

zepdrix (zepdrix):

\[\large \frac{-10}{5-7i}\] `How do you know to... ` Yes, it's because we have an imaginary, or complex rather, value in the denominator. So we want to get that out of there. So what is the conjugate of 5-7i? :D

OpenStudy (anonymous):

Thanks I want to make sure I know this info for tomorrow! 5+7i/5t7i

OpenStudy (anonymous):

Safe to say when it states simplify the complex number and it's in the denominator I use conjugate correct?

zepdrix (zepdrix):

Yah, we don't want to leave complex numbers in the denominator c: The same thing comes up when you deal with square roots. We don't leave `irrational` numbers in the denominator if we can avoid it. Example:\[\large \frac{1}{\sqrt 2} \qquad = \qquad \frac{2}{\sqrt 2}\] We remove the irrational number \(\large \sqrt2\) from the denominator using some multiplication. Blah I got off topic a bit there :D lol

zepdrix (zepdrix):

Woops, I wrote that backwards T.T\[\large \frac{1}{\sqrt 2} \qquad = \qquad \frac{\sqrt2}{ 2}\]

OpenStudy (anonymous):

do you mind breaking down the hook part of the square root sign and how I could make it into a fraction we could use example \[\sqrt[3]{8x^3}\]

zepdrix (zepdrix):

We can rewrite a root as a `rational expression` by using a fraction in our exponent, as you mentioned. Example:\[\large \sqrt {x^3} \qquad = \qquad x^{3/2}\] The numerator of our fraction represents the POWER on x. The denominator represents the degree of the root. When no number is shown in front of the root, it signifies the `square` root, or second root. As I'm sure you're aware.

zepdrix (zepdrix):

With the example you gave, we don't really need to use this idea. We first recognize that 8 is a `perfect cube`. 8=2*2*2. \[\large \sqrt[3]{8x^3} \qquad = \qquad \sqrt[3]{2^3x^3}\qquad = \qquad \sqrt[3]{(2x)^3}\]

OpenStudy (anonymous):

Right what if instead of 8 place 7 would that work in making a rational expression

zepdrix (zepdrix):

Sure, let's try that.\[\large \sqrt[3]{7x^3} \qquad = \qquad \left(7x^3\right)^{1/3}\] We can rewrite the 3rd root as 1/3. The one power (no change), and a 3 in the denominator letting us know it's a third root. From here, we need to remember our rules for exponents.

zepdrix (zepdrix):

One important rule is: We have to apply the exponent to EACH VALUE inside.\[\large \left(7x^3\right)^{1/3} \qquad = \qquad 7^{1/3}(x^3)^{1/3}\]

zepdrix (zepdrix):

Another rule: When you have a term being raised to a power, raised to a power, you `multiply` the exponents. So we simply multiply the 3 and 1/3.\[\large 7^{1/3}(x^3)^{1/3} \qquad = \qquad 7^{1/3}x^{3/3}\qquad = \qquad 7^{1/3}x\]

OpenStudy (anonymous):

Do you mind showing me a few more examples on how to make rational expressions when given them in square root form

zepdrix (zepdrix):

Let's try this one. \[\large \sqrt{4x^3}\]

OpenStudy (anonymous):

so it's (4x)^2/3?

zepdrix (zepdrix):

Maybe it will make more sense if we don't rationalize the root right away. I have a feeling that step is confusing you. Let's try simplifying first.\[\large \sqrt{4x^3} \qquad = \qquad \sqrt4 \cdot \sqrt{x^3}\]All I did was split up the roots so far.

OpenStudy (anonymous):

Okay

zepdrix (zepdrix):

Sqrt4=?

OpenStudy (anonymous):

Got it so far 2

zepdrix (zepdrix):

\[\large \sqrt4 \cdot \sqrt{x^3} \qquad = \qquad 2 \sqrt{x^3}\] Ok good. So from here, we'll apply this idea of writing it as a rational expression. It's a second root (our denominator will be a 2) and a 3 power (our numerator will be a 3).

OpenStudy (anonymous):



OpenStudy (anonymous):

so is it x^2/3?

zepdrix (zepdrix):

Where did your 2 in front go? :D Don't forget about that 2!

OpenStudy (anonymous):

2x*^2/3

OpenStudy (anonymous):

2x^2/3

zepdrix (zepdrix):

3 was our power. It should be on TOP of the fraction.

zepdrix (zepdrix):

the 2 was the degree of the root, DENOMINATOR :D

zepdrix (zepdrix):

\[\large 2 \sqrt{x^3} \qquad = \qquad 2x^{3/2}\]

OpenStudy (anonymous):

when its x^n inside the square root its the denominator?